Find the value of each angle of a triangle given that:
1st angle = x
2nd angle = 4x
3rd angle = (x + 4x) - 45
Armed with this information, find each of the requested angle measures.
1
Expert's answer
2011-11-18T08:30:24-0500
We know that the sum of all angles in a triangle is 180: 1st angle+2nd angle + 3rd angle = 180
X+4x+ x+4x-45=180
10x=225
1st angle:& 22.5
2nd angle: 22.5*4=90
3rd angle:& (22.5 + 90)-45=67.5